the lines described by y = (8a-8)x and y = 1/5x are perpendicular. What is the value of a ?



Answer by stanbon(75887)

the lines described by y = (8a-8)x and y = 1/5x are perpendicular. What is the value of a ?
----
The product of the slopes must be -1.
(8a-8)(1/5) = -1
======
8a-8 = -5
8a = 3
a = 3/8
----------------
Cheers,
Stan H.
---------------